# FLINT: Learning-based Flow Estimation and Temporal Interpolation for Scientific Ensemble Visualization

### v1

@@ III-D Comparison between FLINT and RIFE methods
Figure 5: Droplets ensemble: FLINT flow field estimation and temporal density interpolation during inference—at the timesteps with an asterisk (*)—3$\times$ interpolation. From top to bottom, the rows show GT density, FLINT interpolated density, difference to GT density (magnified by ×$\times 5$), RIFE interpolation, difference to GT density (magnified by ×$\times 5$), FLINT flow estimation in HSV (see bottom right), and RIFE flow estimation.
IV Study Setup
In this section, we describe the training setup, provide an overview of the datasets used in our experiments and discuss the evaluation methods employed to assess the results obtained by FLINT.
IV-A Training
We apply the standard prepossessing step of normalization to [0, 1] when necessary before the start of the training. FLINT is optimized using AdamW [50] — an adaptive gradient descent method with weight decay used in back-propagation algorithms for training feed-forward neural networks that combines the benefits from both the Adam optimizer [51] and $L_{2}$-regularization. We employ early stopping with a patience parameter of 30, which is equivalent to regularization [52] and helps to prevent overfitting on the training data. We use an experimentally determined learning rate of $6\times 10^{-4}$ for the 2D case and $1\times 10^{-4}$ for the 3D case respectively with a cosine annealing scheduler that gradually decreases the learning rate to $6\times 10^{-6}$ and $1\times 10^{-6}$ respectively by the end of the training. We train FLINT with mini-batches of size 32 for both {old}the LBS and Droplets ensembles 2D ensemble datasets and mini-batches of size 2 for 3D datasets used in our study (see Sec. IV-B). We split the set of all available data into training, validation (for monitoring training progress), and test subsets. Note that training and test data are obtained by subsampling the original dataset, which is assumed to provide the GT scalar fields. Then FLINT performs interpolation at missing timesteps.
IV-B Datasets
We consider four scientific datasets in our study.
Nyx (flow-supervised). The fourth dataset is a 3D+time ensemble based on the compressible cosmological hydrodynamics simulation Nyx, developed by Lawrence Berkeley National Laboratory [54]. We consider an ensemble comprising 18 members, each consisting of a maximum of 1600 timesteps with a spatial resolution of $128\times 128\times 128$. It contains density and the $x$, $y$, $z$ components of velocity. Akin to InSituNet [55], we vary three parameters for ensemble generation: the total matter density ($\Omega_{m}\in[0.12,0.155]$), the total density of baryons ($\Omega_{b}\in[0.0215,0.0235]$), and the Hubble constant ($h\in[0.55,0.7]$). We randomly sample a training subset of 500 timesteps and utilize different ensemble members for training, validation, and testing.
To visualize a simulation ensemble field, we use the Turbo colormap [56]. We visualize the 2D flow field via arrow glyphs and 3D flow via volume rendering with a transfer function that indicates flow direction. Examples of the LBS and Droplets ensemble fields, as well as 5Jets and Nyx transfer functions, are in the supplemental material.
IV-C Evaluation
{old}
@@ IV-C Evaluation
Figure 6: 5Jets: FLINT flow field estimation and temporal density interpolation during inference, 20$\times$. From top to bottom, the rows show GT density, FLINT interpolated density, CoordNet interpolation, GT flow, and FLINT flow estimation. 3D rendering was used for the density and flow field visualization ( colors representing $x$, $y$, and $z$ flow directions respectively), transfer functions are in the supplementary material.
@@
V Qualitative Results
V-A LBS: Flow Field Available for Some Members
[…]
In general, going from an interpolation rate of 6 (Fig. 4a) to 18 (Fig. 4b), the density interpolation remains quite accurate (differences are only visible due to significant {old}scalingcontrast enhancement). Naturally, there are more errors in the flow predictions of both FLINT and RIFE when increasing the interpolation rate, but FLINT is still able to maintain comparably high accuracy. Further examples including different interpolation rates can be found in the supplemental material.
Additionally, we perform a streamline analysis to further assess the accuracy of the flow estimated by FLINT. Streamlines were generated to visualize the flow field, with results shown in the last rows, and compared against GT streamlines in the second-to-last rows in Fig. 4. It can be seen that the streamline patterns produced by FLINT align closely with the GT for an interpolation rate of 6 (Fig. 4a), indicating a high level of accuracy in the flow field prediction. Even at an increased interpolation rate of 18 (Fig. 4b), the streamline patterns closely reflect the GT, although some small deviations become noticeable (e.g., for $t=4614$ and $t=4617$). These findings further validate the robustness of FLINT in capturing complex flow dynamics across varying interpolation rates, demonstrating the utility of reconstructing this otherwise missing flow information for analyzing flow around a cylinder.
V-B Droplets: No GT Flow Field Available
The effectiveness of FLINT in interpolating the luminance field—captured by cameras during an experiment—based on the estimated optical flow, even in the absence of GT flow information, is evident from Fig. 5. The results demonstrate camera image interpolation by FLINT of high quality for the Droplets ensemble, in the second row. We compare FLINT’s temporal interpolation performance to the one achieved with RIFE in Fig. 5, fourth row. RIFE’s interpolation exhibits slightly higher variation compared to the GT, as can be seen from the difference plots, in the third and fifth rows.
V-C 5Jets: Density & Flow Available for Some Timesteps
@@ V-D Nyx: Density & Flow Available for Some Timesteps
Fourth, we consider the scenario where GT density and velocity fields are available for some members of the entire 3D ensemble. As Fig. 7 illustrates, FLINT achieves accurate performance in terms of both density field interpolation and flow field estimation. Visually, the difference between the renderings of the reconstructed density field and its GT is minimal. Moreover, even at a relatively high interpolation rate of 5×5$\times$5 ×, FLINT effectively learns a flow field that structurally resembles the GT flow. When comparing our FLINT method to STSR-INR for temporal density interpolation (top three rows), it is evident that the interpolation of the density field is more accurate with FLINT. For example, at $t=503$, STSR-INR shows a different structure and less dark matter density compared to FLINT, which preserves density. This trend is consistently observed across all shown timesteps in Fig. 7. For this comparison, we used the same model size for both FLINT and STSR-INR.
@@
Figure 8: Comparison of FLINT, RIFE, and linear interpolation at various interpolation rates. For PSNR plots for the Droplets dataset and LPIPS scores for the LBS dataset, see Figures 13 and 14 of the supplementary material, respectively.
@@ VI-A Comparison Against Baselines
In the case of the Droplets ensemble, FLINT significantly outperforms linear interpolation for both LPIPS and PSNR metrics, confirming our expectations, as shown in Fig. 8c and supplemental material. This superior performance can be attributed to FLINT’s capacity to learn robust optical flow, which in turn enhances its ability to interpolate and reconstruct complex density fields. FLINT consistently performs better than RIFE as well, across all presented interpolation rates, for both PSNR and LPIPS metrics. This performance reaffirms that our proposed model is well-suited for the challenging task of reconstructing density fields within scientific ensembles. These findings underscore the potential of FLINT in spatiotemporal data interpolation tasks, with application for scientific visualization and data analysis.
@@
We conducted an inference time comparison of our model, FLINT, against RIFE, CoordNet, and STSR-INR across all datasets. Specifically, we compared FLINT with RIFE on the 2D ensembles (LBS and Droplets) and with CoordNet and STSR-INR on the 3D datasets (5Jets and Nyx). For each evaluation, we measured the inference time over 3K timesteps for the 2D ensembles and 300 timesteps for the 3D ensembles from the test set. Our findings indicate that FLINT achieves faster inference time compared to RIFE, with an average of 0.025 seconds per timestep on the 2D datasets, as opposed to RIFE’s 0.035 seconds. On the 3D datasets, FLINT significantly outperforms both CoordNet and STSR-INR in inference speed, even achieving interactive rates with an average time of 0.2 seconds per timestep for FLINT. In comparison, it takes 2.1 seconds for CoordNet and 1.5 seconds for STSR-INR. The faster inference time of FLINT compared to INR-based approaches like CoordNet and STSR-INR can be attributed to the inherent efficiency of convolutional neural networks (CNNs), which excel in parallel processing and optimized memory access, while INR-based methods require solving implicit functions at each point, adding computational complexity.
@@ VI-B Ablation Studies
TABLE III: Ablation of FLINT (flow-supervised)
@@
TABLE IV: Ablation of FLINT, Droplets, 2$\times$ (flow-unsupervised)
@@
The ablation studies conducted on FLINT provide valuable insights into the impact of different loss components on the overall performance. In Table IV, which presents the results for the LBS, 5Jets, and Nyx datasets with available GT flow fields for the model training, four variants of FLINT are compared: “FLINT no flow”, “FLINT no rec”, “FLINT w/o s-t”, and “FLINT”. “FLINT no flow” refers to FLINT without the inclusion of the flow-related loss component. As expected, this variant demonstrates considerably less accurate physical flow, measured by EPE metric, despite a relatively high density interpolation score, measured by PSNR {old}score metric. “FLINT no rec” refers to FLINT without the reconstruction-related loss component. In this case, the {old}physical flow learned by the model is relatively good, however, the density interpolation suffers due to the absence of its related loss component. “FLINT w/o s-t” refers to FLINT {old}without the student-teacher architecture where the teacher block (see Fig. 2) and the teacher-related terms in the loss functions are removed. This scenario leads to slightly less accurate interpolation and flow estimation due to the fact that the model does not receive GT densities during the training. Finally, “FLINT” represents the complete FLINT model with all loss components. It achieves the highest scores across all metrics, including PSNR and EPE, outperforming {old}both the {old}baseline RIFE method and the “FLINT no flow” variant all other variants. This demonstrates the significance of incorporating the proposed loss components in achieving superior interpolation results.
Similar ablation studies were performed for the Droplets ensemble, for which no GT flow fields are available, as shown in Table IV. Multiple variants of FLINT were evaluated, including “FLINT no rec”, “FLINT no dis” (without distillation), “FLINT no photo” (without photometric loss), “FLINT no reg” (without regularization), ablation of combinations of these losses, and “FLINT w/o s-t”. Each variant is compared to the complete FLINT method. The results indicate that each loss component contributes to the overall performance of FLINT. The complete FLINT model outperforms all the variants in terms of PSNR and LPIPS scores. This showcases the importance of reconstruction, distillation, photometric, and regularization loss components in achieving the best possible reconstructions for the Droplets ensemble.
@@ VI-C Parameter Studies
Our proposed FLINT method involves several hyperparameters, and we performed extensive parameter studies to understand their impact on achieving optimal results. Table V displays the outcomes of these studies regarding 2D+time and 3D+time datasets.
TABLE V: Hyperparameter search for FLINT
@@
The parameter search investigates various configurations of FLINT , each labeled with a specific identifier. To determine the model’s optimal architecture for our task, we conducted a series of experiments, adjusting the model’s capacity by varying its width and depth. For example, “FLINT 128” represents a configuration with 128 channels in all convolutional layers across all blocks. Through hyperparameter optimization, we found that the best-performing model, named “FLINT” or “FLINT 4 Blocks”, increases both the channel capacity and the number of blocks. This configuration, with 4 blocks and channel counts in the convolutional layers ranging from 256 to 128, is designed to capture more intricate and detailed features within the data. Another configuration is “FLINT smooth”, where a smoothness loss component was incorporated into the training process for Droplets, however, not yielding the best results. Additionally, “FLINT stride = 1” represents an architecture where each block uses convolutional and deconvolutional layers with a stride of one, instead of varying strides. Furthermore, we explored two different loss functions for the interpolation part of the model. The configuration “FLINT Lapl” indicates the use of the Laplacian loss, which measures the $L_{1}$-loss between two Laplacian pyramid representations (pyramidal level is 5) of the reconstructed density field and GT density field. The configuration “FLINT” employs the simple $L_{1}$-distance as the loss function, as described in Sec. III-C. In addition, we varied the number of blocks of FLINT (Conv Block in Fig. 2), ranging from three to six. As can be seen from Table V, the most optimal configuration is the one with four blocks (“FLINT 4 Blocks”). Models with fewer than four blocks lack sufficient capacity, while those with more begin to overfit.
The results in Table V demonstrate optimized FLINT’s superior performance over other configurations, with the highest scores in PSNR and EPE metrics. This underscores the importance of channel capacity, number of blocks, and optimized loss functions in capturing the intricate dynamics of fluid ensembles, enabling FLINT’s best density interpolation and flow estimation capabilities.

### v2

@@ IV-C Evaluation
Figure 5: Droplets ensemble: FLINT flow field estimation and temporal density interpolation during inference—at the timesteps with an asterisk (*)—3$\times$ interpolation. From top to bottom, the rows show GT density, FLINT interpolated density, difference to GT density (magnified by ×$\times 5$), RIFE interpolated density, difference to GT density (magnified by ×$\times 5$), FLINT flow estimation in HSV (see bottom right), RIFE flow estimation in HSV, flow glyphs for FLINT, and flow glyphs for RIFE.
V Qualitative Results
V-A LBS: Flow Field Available for Some Members
[…]
Additionally, we perform a pathline analysis to further assess the accuracy of the flow estimated by FLINT. Pathlines were generated to visualize the flow field. Results for FLINT are shown in the last row and GT pathlines in the last but one row in Fig. 4. It can be seen that the pathline patterns produced by FLINT align closely with the GT for an interpolation rate of 6 (Fig. 4a), indicating a high level of accuracy in the flow field estimation. Even at an increased interpolation rate of 18 (Fig. 4b), the pathline patterns closely reflect the GT, although some small deviations become noticeable (e.g., for $t=4608$, $t=4614$, and $t=4617$). These findings further validate the robustness of FLINT in capturing complex flow dynamics across varying interpolation rates, demonstrating the utility of reconstructing this otherwise missing flow information for analyzing flow around a cylinder.
@@ V-A LBS: Flow Field Available for Some Members
Figure 6: 5Jets: FLINT flow field estimation and temporal density interpolation during inference, 20$\times$. From top to bottom, the rows show GT density, FLINT interpolated density, CoordNet interpolated density, GT flow, and FLINT estimated flow. 3D rendering was used for the density and flow field visualization ( colors representing $x$, $y$, and $z$ flow directions respectively), transfer functions are in the supplementary material.
V-B Droplets: No GT Flow Field Available
The effectiveness of FLINT in interpolating the luminance field—captured by cameras during an experiment—based on the estimated optical flow, even in the absence of GT flow information, is evident from Fig. 5. The results demonstrate camera image interpolation by FLINT of high quality for the Droplets ensemble, in the second row. We compare FLINT’s temporal interpolation performance to the one achieved with RIFE in Fig. 5, fourth row. RIFE’s interpolation exhibits slightly higher variation compared to the GT, as can be seen from the difference plots, in the third and fifth rows. It is noteworthy that this concerns experimental data that naturally contains noise. For instance, in the top row of Fig. 5, the noise is visible as small, scattered inconsistencies in the background throughout the luminance field, giving it a slightly grainy appearance. Additionally, shadows can be observed on the right side of the images, where darker regions obscure portions of the luminance field. Despite this, FLINT demonstrates robust performance, effectively handling these imperfections and delivering high-quality interpolation results.
V-C 5Jets: Density & Flow Available for Some Timesteps
@@ V-D Nyx: Density & Flow Available for Some Timesteps
Fourth, we consider the scenario where GT density and velocity fields are available for some members of the entire 3D ensemble. As Fig. 7 illustrates, FLINT achieves accurate performance in terms of both density field interpolation and flow field estimation. Visually, the difference between the renderings of the reconstructed density field and its GT is minimal. Moreover, even at a relatively high interpolation rate of 5×5$\times$5 ×, FLINT effectively learns a flow field that structurally resembles the GT flow. When comparing our FLINT method to STSR-INR for temporal density interpolation (top three rows), it is evident that the interpolation of the density field is more accurate with FLINT. For example, at $t=503$, STSR-INR shows a different structure and less dark matter density compared to FLINT, which preserves density. This trend is consistently observed across all shown timesteps in Fig. 7. For this comparison, we used the same model size for both FLINT and STSR-INR. By default, Implicit Neural Representation (INR)-based models like CoordNet and STSR-INR have smaller model sizes compared to FLINT due to their simpler architecture, where data is represented as a continuous function parameterized by a compact set of weights. In contrast, FLINT employs a more complex design with multiple convolutional blocks, enabling it to achieve accurate flow estimation and scalar field interpolation. To ensure a fair comparison below, we configured STSR-INR to have the same model size as FLINT, allowing us to directly assess their performance under comparable conditions.
@@
V-E Nyx: Domain Expert Evaluation
A domain expert—an assistant professor in astronomy with a research focus on cosmological simulations and observational data evaluation—provided insights into FLINT’s utility for analyzing scalar fields and velocity information in the Nyx dataset. He highlighted that FLINT’s ability to estimate velocity fields alongside density fields offers significant advantages. The expert remarked, “compared to other reconstruction methods, with FLINT you get information not only on the density in Nyx simulations but also on the velocity”. This dual capability is particularly important in cosmology, where velocity data derived from “redshift” measurements are critical for estimating distances and analyzing spatial relationships between objects.
The expert stated that “having a way to estimate accurately the peculiar velocities would help when comparing with observations”. He further underlined the relevance of velocity information in creating “lightcones”, which combine simulation timesteps to mimic observations of the universe at different epochs. These lightcones are crucial for studying phenomena such as the intergalactic medium and the 3D distribution of galaxies, where “the distance information comes really from a velocity shift”. He emphasized that “having a way to estimate accurately the velocities would help when comparing with observations”. By accurately reconstructing both density and velocity fields, FLINT enhances the fidelity of simulation outputs aligned with observational data, bridging gaps between simulations and real-world measurements, and enabling a deeper understanding of cosmological dynamics.
@@ V-E Nyx: Domain Expert Evaluation
Figure 8: Comparison of FLINT, RIFE, and linear interpolation at various interpolation rates. For PSNR plots for the Droplets dataset and LPIPS scores for the LBS dataset, see Figures 16 and 17 of the supplementary material, respectively.
@@ VI-A Comparison Against Baselines
In the case of the Droplets ensemble, FLINT significantly outperforms linear interpolation for both LPIPS and PSNR metrics, confirming our expectations, as shown in Fig. 8c and supplementary material. This superior performance can be attributed to FLINT’s capacity to learn robust optical flow, which in turn enhances its ability to interpolate and reconstruct complex density fields. FLINT consistently performs better than RIFE as well, across all presented interpolation rates, for both PSNR and LPIPS metrics. This performance reaffirms that our proposed model is well-suited for the challenging task of reconstructing density fields within scientific ensembles. These findings underscore the potential of FLINT in spatiotemporal data interpolation tasks, with application for scientific visualization and data analysis.
@@
Evaluation on Additional Scalar Field. In addition to evaluating FLINT for density and flow fields, we extended our analysis to another scalar field to demonstrate its versatility. Specifically, we evaluated FLINT’s performance on the temperature field from the Nyx simulation ensemble. We examined FLINT’s ability to interpolate the temperature field and estimate the corresponding flow fields. Qualitative results are shown in Fig. 19 of the supplementary material, and quantitative evaluations are provided in Table III. The temperature with flow results indicate that the temperature interpolation performance remains robust across different interpolation rates, although the PSNR values are slightly lower compared to density results. This difference reflects the increased complexity of interpolating temperature fields due to their distinct data characteristics. Similarly, the flow results show lower EPE scores. This is expected, as estimating flow from temperature is inherently more challenging in comparison to density due to its weaker correlation with movement or physical flow dynamics.
Our findings align with observations by Gu et al. [16], where the performance of vector field reconstruction can vary depending on the input scalar field. Similarly, FLINT’s performance showed variations when applied to the temperature field, reflecting the challenges associated with different data characteristics.
These results further demonstrate FLINT’s adaptability across diverse scalar fields while highlighting the influence of field-specific dynamics on reconstruction accuracy.
TABLE III: Comparison of different fields, Nyx
Field 3$\times$ 5$\times$ 8$\times$
Density + Flow 53.17 0.0310 52.31 0.0310 49.39 0.0309
Temperature + Flow 49.32 0.0446 46.91 0.0451 44.43 0.0467
Impact of Noise. To evaluate FLINT’s robustness to noise, we introduced random Gaussian noise into the 3D Nyx simulation ensemble. Even with the added noise in the GT density, FLINT achieved reliable results, maintaining an average PSNR of 46.36 dB and an EPE of 0.0356 at a 5×5$\times$5 × interpolation rate. In comparison, without noise, the model achieved a higher PSNR of 52.31 dB and a slightly lower EPE of 0.0310, demonstrating a modest decline in performance due to noise. This highlights FLINT’s resilience to noisy data and reinforces its applicability to real-world scenarios where imperfections in scientific datasets are common. For further details, please refer to the supplementary material.
Large Interpolation Rates. At large interpolation rates, such as 32×32$\times$32 × (Fig. 13 of the supplementary material), we observe that while density interpolation maintains acceptable quality (average PSNR of 37.71 dB), the accuracy of flow estimation noticeably degrades (average EPE is 0.0479). In comparison, at 8×8$\times$8 × interpolation, the model achieved a PSNR of 49.39 dB and an EPE of 0.0309, while at 16×16$\times$16 ×, the PSNR was 43.42 dB with an EPE of 0.0338. The degradation at a large interpolation rate is expected due to the increased structural differences between flows at widely separated time steps, which make flow estimation inherently more challenging. Nevertheless, FLINT delivers satisfactory results for scalar field interpolation even at these demanding rates. Qualitatively, while the main structure, such as the distribution of dark matter, is preserved and remains visually consistent with the GT, finer details and intricate patterns become less accurate. Further details are provided in the supplementary material.
@@
Inference time comparison. We conducted an inference time comparison of our model, FLINT, against RIFE, CoordNet, and STSR-INR across all datasets. Specifically, we compared FLINT with RIFE on the 2D ensembles (LBS and Droplets) and with CoordNet and STSR-INR on the 3D datasets (5Jets and Nyx). For each evaluation, we measured the inference time over 3K timesteps for the 2D ensembles and 300 timesteps for the 3D ensembles from the test set. Our findings indicate that FLINT achieves faster inference time compared to RIFE, with an average of 0.025 seconds per timestep on the 2D datasets, as opposed to RIFE’s 0.035 seconds. On the 3D datasets, FLINT significantly outperforms both CoordNet and STSR-INR in inference speed, even achieving interactive rates with an average time of 0.2 seconds per timestep for FLINT. In comparison, it takes 2.1 seconds for CoordNet and 1.5 seconds for STSR-INR. The faster inference time of FLINT compared to INR-based approaches like CoordNet and STSR-INR can be attributed to the inherent efficiency of convolutional neural networks (CNNs), which excel in parallel processing and optimized memory access, while INR-based methods require solving implicit functions at each point, adding computational complexity.
@@ VI-B Ablation Studies
TABLE IV: Ablation of FLINT (flow-supervised)
@@
TABLE V: Ablation of FLINT, Droplets, 2$\times$ (flow-unsupervised)
@@
The ablation studies conducted on FLINT provide valuable insights into the impact of different loss components on the overall performance. In Table V, which presents the results for the LBS, 5Jets, and Nyx datasets with available GT flow fields for the model training, four variants of FLINT are compared: “FLINT no flow”, “FLINT no rec”, “FLINT w/o s-t”, and “FLINT”. “FLINT no flow” refers to FLINT without the inclusion of the flow-related loss component. As expected, this variant demonstrates considerably less accurate physical flow, measured by EPE metric, despite a relatively high density interpolation score, measured by PSNR {old}score metric. “FLINT no rec” refers to FLINT without the reconstruction-related loss component. In this case, the {old}physical flow learned by the model is relatively good, however, the density interpolation suffers due to the absence of its related loss component. “FLINT w/o s-t” refers to FLINT {old}without the student-teacher architecture where the teacher block (see Fig. 2) and the teacher-related terms in the loss functions are removed. This scenario leads to slightly less accurate interpolation and flow estimation due to the fact that the model does not receive GT densities during the training. Finally, “FLINT” represents the complete FLINT model with all loss components. It achieves the highest scores across all metrics, including PSNR and EPE, outperforming {old}both the {old}baseline RIFE method and the “FLINT no flow” variant all other variants. This demonstrates the significance of incorporating the proposed loss components in achieving superior interpolation results.
Similar ablation studies were performed for the Droplets ensemble, for which no GT flow fields are available, as shown in Table V. Multiple variants of FLINT were evaluated, including “FLINT no rec”, “FLINT no dis” (without distillation), “FLINT no photo” (without photometric loss), “FLINT no reg” (without regularization), ablation of combinations of these losses, and “FLINT w/o s-t”. Each variant is compared to the complete FLINT method. The results indicate that each loss component contributes to the overall performance of FLINT. The complete FLINT model outperforms all the variants in terms of PSNR and LPIPS scores. This showcases the importance of reconstruction, distillation, photometric, and regularization loss components in achieving the best possible reconstructions for the Droplets ensemble.
@@ VI-C Parameter Studies
Our proposed FLINT method involves several hyperparameters, and we performed extensive parameter studies to understand their impact on achieving optimal results. Table VI displays the outcomes of these studies regarding 2D+time and 3D+time datasets.
TABLE VI: Hyperparameter search for FLINT
@@
FLINT two-stage 56.87 0.0069 45.78 0.7409 52.07 0.0310 40.95 0.0108
The parameter search investigates various configurations of FLINT , each labeled with a specific identifier. To determine the model’s optimal architecture for our task, we conducted a series of experiments, adjusting the model’s capacity by varying its width and depth. For example, “FLINT 128” represents a configuration with 128 channels in all convolutional layers across all blocks. Through hyperparameter optimization, we found that the best-performing model, named “FLINT” or “FLINT 4 Blocks”, increases both the channel capacity and the number of blocks. This configuration, with 4 blocks and channel counts in the convolutional layers ranging from 256 to 128, is designed to capture more intricate and detailed features within the data. Another configuration is “FLINT smooth”, where a smoothness loss component was incorporated into the training process for Droplets, however, not yielding the best results. Additionally, “FLINT stride = 1” represents an architecture where each block uses convolutional and deconvolutional layers with a stride of one, instead of varying strides. Furthermore, we explored two different loss functions for the interpolation part of the model. The configuration “FLINT Lapl” indicates the use of the Laplacian loss, which measures the $L_{1}$-loss between two Laplacian pyramid representations (pyramidal level is 5) of the reconstructed density field and GT density field. The configuration “FLINT” employs the simple $L_{1}$-distance as the loss function, as described in Sec. III-C. In addition, we varied the number of blocks of FLINT (Conv Block in Fig. 2), ranging from three to six. As can be seen from Table VI, the most optimal configuration is the one with four blocks (“FLINT 4 Blocks”). Models with fewer than four blocks lack sufficient capacity, while those with more begin to overfit. The configuration “FLINT two-stage” employs a two-stage optimization approach where a teacher model is trained first, followed by a student network trained to align with the teacher’s outputs. In contrast, “FLINT” achieves slightly better performance overall while reducing training time by nearly half (12 vs. 20 hours), demonstrating its efficiency.
The results in Table VI demonstrate optimized FLINT’s superior performance over other configurations, with the highest scores in PSNR and EPE metrics. This underscores the importance of channel capacity, number of blocks, and optimized loss functions in capturing the intricate dynamics of fluid ensembles, enabling FLINT’s best density interpolation and flow estimation capabilities.